displacement-time equation of a particle executing shm is x=asin(wt+π/6) time taken by particle to go directly from x=-a/2 to x=+a/2 is

Grade:12

2 Answers

Khimraj
5 years ago
initial position of particle
at t=0
x = asin(π/6) = a/2.
amplitude of particle = a
so time taken by particle to move from -a/2 to a/2 will be T/4.
shreesharma
2 years ago
x=a sin(wt+pi/6), therefore it’s initial position wil a/2 moving towards mean position.Time taken by the particle to go from a/2 to -a/2 will be equal to the time taken by the particle to go from -a/2 to a/2.
x=a sin(wt+pi/6)
put x=-a/2
-a/2=a/2 sin(wt+pi/6)
therefore, wt+pi/6=-pi/6
wt=-pi/3
t=-pi/3w hence time can never be in negative therefore t=pi/3w